OpenStudy (anonymous):

find solutions for cos(2x)=4 cos x-2

OpenStudy (anonymous):

the 1st one is 1/2d to the 2 power fin to the 2 e2 u

OpenStudy (anonymous):

could you break it down?

OpenStudy (anonymous):

x=12d2fin2e2U

OpenStudy (anonymous):

x=1/2d2fin2e2U

OpenStudy (anonymous):

sorry i need solutions from 0 to 2pi. so in radians

OpenStudy (anonymous):

2cos^2 x -1 = 4cos x - 2. Call x = cos x 2x^2 - 4x + 1 = 0 . quadratic equation with 2 real roots x1 = 1 + V2/2 (rejected as > 1), and x2 = 2 - V2/2 = 0.293 cos x = x2 = 0.293 = cos A x = A or x = -A. (check)